### v2

@@ 3.2 Learning-Augmented Densest At-Most-$k$ Subgraph
To make the brute-force computation more efficient, we incorporate a pruning procedure. Specifically, we iteratively remove nodes with minimum degree from the graph until only $k$ nodes remain. Let the resulting graph be denoted by $H$, and let $d^{\prime}=\textup{d}(H)$. Starting from $H$, we again iteratively remove nodes of minimum degree and record the maximum density attained by any intermediate subgraph during this process. Since $H$ is a subgraph with at most $k$ nodes, $d^{\prime}$ serves as a lower bound on the density of the densest subgraph with at most $k$ nodes.
@@
$>\frac{e(H^{\star})-d^{\star}}{|H^{\star}|-1}$
@@
This is a contradiction, since $H^{\star}\setminus\{v\}$ is also a feasible solution and achieves strictly higher density than $H^{\star}$. Therefore, the removed nodes cannot be part of an optimal solution. This pruning rule allows us to significantly speed up our exact algorithms.
Generating a labeled dataset on 800 graphs (which involves finding the ground truth densest at-most-$k$ subgraph) took 75 minutes on 8 threads in Julia. Ignoring some graphs that are too big in our experiment, we obtain 592 graphs for training a predictor and 149 graphs for testing. Our learning-augmented algorithm finishes in less than a second on 149 testing input graphs in Python. Similar to the densest subgraph experiment, we train a random forest and treat it as a predictor. In this experiment, we use $\epsilon=0.1$ as the random forest seems to perform better than in the densest subgraph case (see Table 4).
Figure 6: An example of densest at-most-15 densest subgraph in Twitch’s ego-networks. Note that the at most constraint avoid a large densest subgraph output.
Class Precision Recall F1-score Support
0 0.88 0.87 0.88 2522
1 0.81 0.81 0.81 1636
Accuracy 0.85 1458
Table 3: Classification report of a random forest to predict whether a node belongs to the densest at-most-15 subgraph using only degree, average neighbor degree, and graph size as features. 0 means not being in the densest subgraph. 1 means being in the densest subgraph.
We compare the density obtained by our algorithm in Section 2.2, the optimal solution, and the density of the predictor’s solution (i.e., the nodes that are predicted to be in the densest at-most-15 subgraph by the random forest). We emphasize that the predictor’s solution might have more than $k$ nodes and may not be a feasible solution. See Figure 7.
@@
Average Approximation Ratio 0.87
Median Approximation Ratio 0.95
High Quality Solutions ($\geq 0.95$) 72 / 149 (48.64%)
Table 4: Approximation factors of Algorithm 2 for the densest at-most-$k$ subgraph problem ($k=15$, $\epsilon=0.1$).
We observe that the learning-augmented solution often achieves higher density than the predictor’s output (which might not be feasible when its size exceeds $k$). We also compare our algorithm’s output density vs the optimal density found by brute-force. As seen in Table 5, the average and median approximation factors of our algorithms in the testing set is 0.87 and 0.95 respectively.
@@
Average Approximation Ratio 0.886
Median Approximation Ratio 0.945
High Quality Solutions ($\geq 0.95$) 27 / 55 (51.64%)
Table 5: Approximation of Algorithm 2 with $\epsilon=0.1$ for the densest at-most-$k$ subgraph problem.
Figure 7: Pairwise comparison among our algorithm’s solution (Section 2.2), the optimal solution, and the density of the predicted set of nodes for the densest at-most-$k$ subgraph problem.
Figure 8: Approximation factors of Algorithm 2 on testing graphs for the densest at-most-$k$ subgraph problem.
@@ 4 Conclusion
In this paper, we presented several simple and practical learning-augmented linear-time algorithms for computing the densest subgraph and its variants, including the NP-hard densest at-most-$k$ subgraph. These algorithms offer several desirable properties. They come with provable approximation guarantees and incur minimal computational overhead, running in linear time aside from the cost of executing the predictor. We also demonstrated, both theoretically and empirically, that blindly relying on the predictor can lead to poor approximations or infeasible solutions. When used appropriately, our learning-augmented algorithms can outperform Charikar’s classic greedy algorithm and the predictor’s output alone. In particular, our experiments for the densest at-most-$k$ subgraph show that the learning-augmented approach produces feasible solutions with strong empirical performance, achieving near-optimal density on most test instances. Finally, these methods are highly practical and can be implemented in just a few lines of code.
Appendix A Learning-Augmented Algorithm for Directed Densest Subgraph
Given a directed graph $G=(V,E)$, the density of $(S_{1},S_{2})$, where $S_{1},S_{2}\subseteq V$, is defined as follows.
$\textup{d}(S_{1},S_{2})=\frac{e(S_{1},S_{2})}{\sqrt{|S_{1}|\cdot|S_{2}|}}.$
Here, we define $E(S_{1},S_{2})$ to be the set of edges from a node in $S_{1}$ to a node in $S_{2}$ and $e(S_{1},S_{2})=|E(S_{1},S_{2})|$. Let $S_{1}^{\star},S_{2}^{\star}$ be the optimal solution. In this setting, we assume access to $S_{1},S_{2}\subseteq V$ that are $1-\epsilon$ partial solutions to $S_{1}^{\star}$ and $S_{2}^{\star}$ respectively.
We want to design an algorithm that finds a $1-O(\epsilon)$ approximation to $S_{1}^{\star},S_{2}^{\star}$. We start with the following proposition.
Proposition 2.
Let $S^{\prime}_{i}=S^{\star}_{i}\setminus S_{i}$ for $i=1,2$, then $e(S^{\prime}_{1},S^{\prime}_{2})\leq\sqrt{\epsilon_{1}\epsilon_{2}}\cdot e(S^{\star}_{1},S^{\star}_{2})$.
Proof.
By the optimality of the pair $(S^{\star}_{1},S^{\star}_{2})$, we have $\textup{d}(S^{\prime}_{1},S^{\prime}_{2})\leq\textup{d}(S^{\star}_{1},S^{\star}_{2})$, which gives
$e(S^{\prime}_{1},S^{\prime}_{2})\leq\frac{e(S_{1}^{\star},S_{2}^{\star})}{\sqrt{|S_{1}^{\star}||S_{2}^{\star}|}}\cdot\sqrt{|S^{\prime}_{1}||S^{\prime}_{2}|}\leq\sqrt{\epsilon_{1}\epsilon_{2}}\cdot e(S^{\star}_{1},S^{\star}_{2}).\qed$
For each $v\in V$, we define $w_{S_{1}}(v)$ and $w_{S_{2}}(v)$ as follows
$w_{S_{1}}(v)=\begin{cases}0&\text{ if $v\in S_{1}$ }\\ |\{v\to u\in E:u\in S_{2}\}|&\text{ otherwise }\end{cases}$
and
$w_{S_{2}}(v)=\begin{cases}0&\text{ if $v\in S_{2}$ }\\ |\{u\to v\in E:u\in S_{1}\}|&\text{ otherwise.}\end{cases}$
The algorithm below follows a greedy approach: it first selects $\left\lceil\frac{\epsilon_{1}}{1-\epsilon_{1}}|S_{1}|\right\rceil$ nodes with the highest $w_{S_{1}}$ values and $\left\lceil\frac{\epsilon_{2}}{1-\epsilon_{2}}|S_{2}|\right\rceil$ nodes with the highest $w_{S_{2}}$ values. These nodes are then added to $S_{1}$ and $S_{2}$, respectively, to produce the final output.
Input: Graph $G=(V,E)$, initial sets $S_{1},S_{2}\subseteq V$ that are $1-\epsilon$ partial solutions to $S_{1}^{\star},S_{2}^{\star}$.
1 For each $v\in V(G)$, compute $w_{S_{1}}(v)$ and $w_{S_{2}}(v)$;
2 $U_{1}\leftarrow$ $\left\lceil\frac{\epsilon_{1}}{1-\epsilon_{1}}|S_{1}|\right\rceil$ nodes with the highest $w_{S_{1}}$ values;
3 $U_{2}\leftarrow$ $\left\lceil\frac{\epsilon_{2}}{1-\epsilon_{2}}|S_{2}|\right\rceil$ nodes with the highest $w_{S_{2}}$ values;
4 return $(S_{1}\cup U_{1},S_{2}\cup U_{2})$;
Algorithm 3 Approximate Directed Densest Subgraph with Predictions
Theorem 3.
Let $\epsilon=\max\{\epsilon_{1},\epsilon_{2}\}$. Algorithm 3 yields a
$1-O\left(\epsilon+\frac{1}{|S_{1}^{\star}|}+\frac{1}{|S_{2}^{\star}|}\right)$
approximation to the directed densest subgraph and runs in $O(m)$ time.
Proof.
Set $\overline{S}_{i}:=S_{i}\cap S^{\star}_{i}$ for each $i=1,2,$ we have the following decomposition of $E(S^{\star}_{1},S^{\star}_{2})$
$E(S^{\star}_{1},S^{\star}_{2})=E(\overline{S}_{1},\overline{S}_{2})\sqcup E(\overline{S}_{1},S^{\prime}_{2})\sqcup E(S^{\prime}_{1},\overline{S}_{2})\sqcup E(S^{\prime}_{1},S^{\prime}_{2}).$
The upper bound on $e(S^{\prime}_{1},S^{\prime}_{2})$ and the decomposition of $E(S^{\star}_{1},S^{\star}_{2})$ above imply
$e(\overline{S}_{1},\overline{S}_{2})+e(\overline{S}_{1},S^{\prime}_{2})+e(S^{\prime}_{1},\overline{S}_{2})\geq(1-\sqrt{\epsilon_{1}\epsilon_{2}})e(S^{\star}_{1},S^{\star}_{2}).$
Let $U_{1},U_{2}$ be the sets of $\left\lceil\frac{\epsilon_{1}}{1-\epsilon_{1}}|S_{1}|\right\rceil$ and $\left\lceil\frac{\epsilon_{2}}{1-\epsilon_{2}}|S_{2}|\right\rceil$ nodes with the highest $w_{S_{1}}(\cdot)$ and $w_{S_{2}}(\cdot)$ values respectively.
Observe that $|S_{2}|\geq|\overline{S}_{2}|$ and
$|U_{2}|=\left\lceil\frac{\epsilon_{2}}{1-\epsilon_{2}}|S_{2}|\right\rceil\geq\frac{\epsilon_{2}}{1-\epsilon_{2}}|S_{2}|\geq\frac{1}{1-\epsilon_{2}}|S_{2}^{\prime}|\geq|S_{2}^{\prime}|$
and since we choose $U_{2}$ greedily in terms of indegree from $S_{1}$, we deduce that $e(S_{1},U_{2})\geq e(\overline{S_{1}},S^{\prime}_{2})$. Using a similar argument, we also have $e(U_{1},S_{2})\geq e(S^{\prime}_{1},\overline{S}_{2})$. It is trivial that $e(S_{1},S_{2})\geq e(\overline{S}_{1},\overline{S}_{2})$ as $\overline{S}_{1}\subseteq S_{1},\overline{S}_{2}\subseteq S_{2}$.
The edges in the solution can be decomposed as $E(S_{1}\cup U_{1},S_{2}\cup U_{2})=E(S_{1},S_{2})\sqcup E(S_{1},U_{2})\sqcup E(U_{1},S_{2})\cup E(U_{1},U_{2})$. Therefore,
$e(S_{1}\cup U_{1},S_{2}\cup U_{2})$ $\geq e(\overline{S}_{1},\overline{S}_{2})+e(\overline{S}_{1},S_{2}^{\prime})+e(S_{1}^{\prime},\overline{S}_{2}^{\prime})+0$
$\geq e(S^{\star}_{1},S^{\star}_{2})-e(S^{\prime}_{1},S^{\prime}_{2})\geq(1-\sqrt{\epsilon_{1}\epsilon_{2}})e(S^{\star}_{1},S^{\star}_{2}).$
Then,
$|S_{i}\cup U_{i}|$ $\leq(1+\epsilon_{i})|S^{\star}_{i}|+\left\lceil\frac{\epsilon_{i}}{1-\epsilon_{i}}|S_{i}|\right\rceil$
$\leq(1+\epsilon_{i})|S^{\star}_{i}|+\frac{\epsilon_{i}}{1-\epsilon_{i}}|S^{\star}_{i}|+1$
$<(1+3\epsilon_{i})|S^{\star}_{i}|+1.$
Putting these together, we conclude that
$\textup{d}(S_{1}\cup U_{1},S_{2}\cup U_{2})$ $\geq\frac{1-\sqrt{\epsilon_{1}\epsilon_{2}}}{\sqrt{((1+3\epsilon_{1})|S^{\star}_{1}|+1)((1+3\epsilon_{2})|S^{\star}_{2}|+1)}}\cdot\textup{d}(S_{1}^{\star},S_{2}^{\star})$
$=\left(1-O\!\left(\epsilon+\frac{1}{|S^{\star}_{1}|}+\frac{1}{|S^{\star}_{2}|}\right)\right)\textup{d}(S_{1}^{\star},S_{2}^{\star}).$
Similar to the previous result, it is trivial to implement this algorithm in $O(m)$ time. ∎

### v3

@@ 3.1 Experimental Setup
Specifically, we iteratively remove nodes with minimum degree from the graph until only $k$ nodes remain. Let the resulting graph be denoted by $H$, and let $d^{\prime}=\textup{d}(H)$. Starting from $H$, we again iteratively remove nodes of minimum degree and update the maximum density attained by any intermediate subgraph during this process in $d^{\prime}$ and $H$.
Since $H$ is a subgraph with at most $k$ nodes, $d^{\prime}$ serves as a lower bound on the density of the densest subgraph with at most $k$ nodes.
@@
$>\frac{e(H^{\star})-\textup{d}(H^{\star})}{|H^{\star}|-1}$
@@
This is a contradiction, since $H^{\star}\setminus\{v\}$ is also a feasible solution and achieves strictly higher density than $H^{\star}$. Therefore, the removed nodes cannot be part of an optimal solution. We can then repeat the procedure on the remaining graph until no more nodes can be removed. This pruning rule allows us to significantly speed up our exact algorithms.
We also implemented several other pruning techniques. Let $d^{\prime}$ be the lower bound for the optimal density obtained by the pruning procedure described above. We know that any subset of vertices of size smaller than $2d^{\prime}+1$ cannot have a density higher than $d^{\prime}$; therefore, we can safely ignore all subsets of vertices of size smaller than $2d^{\prime}+1$.
3.2 Twitch Ego Nets
The first dataset we use consists of ego-networks of Twitch users who participated in the partnership program in April 2018, as described in [40]. It contains over 127,000 graphs. Each node represents a user, and each edge represents a friendship.
We restrict our experiment to graph with $n\leq 100$ and use $k=15$.
Figure 3: An example of a densest at-most-15 densest subgraph in the Twitch ego data set.
Since the nodes lack attributes, we follow standard practice in graph learning and use local structural features. Specifically, each node is characterized by its degree, the average degree of its neighbors, and the total number of nodes in its graph. These features can be computed for all nodes in $O(n+m)$ time, matching the linear-time complexity of our algorithm. The label is $1$ if a node belongs to the densest at-most-$15$ subgraph, and 0 otherwise. We obtain the ground truth by brute-force, using the pruning techniques described below.
We use a set of 200 graphs for training and testing, and another 200 graphs for validation. These graphs are chosen uniformly at random. Some instances are dropped because computing the ground truth times out. In total, we retain 179 graphs for training and testing and 173 graphs for validation.
We train our predictor using a random forest classifier with 10 decision trees. For the 179 graphs in the training/testing set, we perform an 80/20 split at the node level: 80% of the nodes are used for training and the remaining 20% for testing. On the testing set, the confusion matrix is shown Table 1.
Predicted 0 Predicted 1
Actual 0 527 83
Actual 1 91 319
Table 1: Confusion matrix for the random forest classifier predicting whether a node belongs to the densest at-most-15 subgraph in the Twitch ego data set.
This suggests a false positive rate of 13.6% and a false negative rate of 22.2%. Hence, we set $\epsilon=0.2$.
We compare the density obtained by our algorithm in Section 2.2, the optimal solution, and the density of the predictor’s solution (i.e., the nodes that are predicted to be in the densest at-most-15 subgraph by the random forest). We emphasize that the predictor’s solution might have more than $k$ nodes and may not be a feasible solution. See Figure 4. The average and median approximation factors of our algorithm in the validation set are 0.81 and 0.94 respectively.
@@ 3.2 Twitch Ego Nets
Average Approximation Ratio 0.806
Median Approximation Ratio 0.943
High Quality Solutions ($\geq 0.95$) 81 / 173 (46.82%)
Table 2: Approximation factors of Algorithm 2 on validation graphs for the densest at-most-$15$ subgraph problem
We also compare the time to compute the exact solution using brute-force with pruning techniques and the learning-augmented algorithm described in Section 2.2 over the validation set. The total runtime of the learning-augmented algorithm over all 173 validation graphs is $2.1$ seconds, compared to $374.52$ seconds for the exact algorithm (a speedup of over $178\times$ on average).
Algorithm Runtime
Learning Augmented algorithm 2.1 s
Exact algorithm (brute force with pruning) 374.52 s
Table 3: Runtime comparison between the learning-augmented algorithm (Section 2.2) and the exact algorithm on validation graphs for the densest at-most-$15$ subgraph problem over 173 graphs in the Twitch ego data set’s validation set.
Figure 4: Pairwise comparison among our algorithm’s solution (Section 2.2), the optimal solution, and the density of the predicted set of nodes for the densest at-most-$15$ subgraph problem in the Twitch ego data set.
Figure 5: Approximation factors of Algorithm 2 on testing graphs for the densest at-most-$15$ subgraph problem in the Twitch ego data set.
3.3 High-energy Physics Citation Network
The arXiv HEP-PH citation graph covers all citations among 34,546 high energy physics phenomenology papers, comprising 421,578 edges. Here, we randomly sample induced subgraphs of size 50 nodes by picking a random node and start a breadth-first search from it until we hit the desired number of nodes. For this experiment, we set $k=10$.
This dataset is significantly more challenging since the graphs are denser everywhere and it is harder to find the exact solution. We had to filter out some hard instances where the exact solution cannot be obtained within a reasonable time frame. In total, the number of graphs for training and testing is 91 and the number of graphs for validation is 96.
Figure 6: An example of the densest at-most-$10$ subgraph in the HEP-PH citation network.
We again train a random forest with 10 decision trees as the classifier using the same features as in the Twitch ego data set. The confusion matrix is as follows:
Predicted 0 Predicted 1
Actual 0 548 47
Actual 1 93 94
Table 4: Confusion matrix of the random forest predictor on the HEP-PH dataset.
From this confusion matrix, we choose $\epsilon=0.497$. Even though the predictor is less accurate than the one for the Twitch ego data set, our learning-augmented algorithm still achieves a respectable average and median approximation factor of 0.81 and 0.93 respectively. Recall that this problem is presumably hard to approximate in general.
Figure 7 shows how important the learning-augmented algorithm is compared to the predictor’s solution. Not only does it makes the solution feasible, but it also significantly improves the density compared to the predictor’s solution.
@@ 3.3 High-energy Physics Citation Network
Average Approximation Ratio 0.8097
Median Approximation Ratio 0.9322
High Quality Solutions ($\geq 0.95$) 46 / 96 (47.92%)
Table 5: Approximation factors of the learning-augmented algorithm on the HEP-PH validation graphs.
Figure 7: Pairwise density comparison on the HEP-PH dataset: augmented algorithm vs. optimal (left) and predictor vs. augmented algorithm (right), for the densest at-most-$k$ subgraph problem.
Figure 8: Distribution of approximation ratios (augmented / optimal) on the HEP-PH validation graphs.
The total runtime of the learning-augmented algorithm over all 96 validation graphs is shown in Table 6.
Algorithm Mean Median Total
Augmented 22.43 ms 0.01 ms 2.15 s
Exact (brute-force with pruning) 14305.8 ms 2223.39 ms 1373.36 s
Table 6: Runtime comparison on the HEP-PH validation graphs (96 instances) between the learning-augmented algorithm and the exact brute-force algorithm.
@@ 4 Conclusion
In this paper, we presented several simple and practical learning-augmented linear-time algorithms for computing the densest subgraph and its variants, including the NP-hard densest at-most-$k$ subgraph. These algorithms offer several desirable properties. They come with provable approximation guarantees and incur minimal computational overhead, running in linear time aside from the cost of executing the predictor. We also demonstrated, both theoretically and empirically, that blindly relying on the predictor can lead to poor approximations or infeasible solutions. When used appropriately, our learning-augmented algorithms provide a practical way to obtain a high-quality solution to the densest at-most-$k$ subgraph problem.